| 
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
Sample Environment
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
 | 
| 
SASE1 Status and SASE3 Readiness
 | 
|   Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  | 
Not logged in | 
  | 
   | 
| 
Message ID: 572
    Entry time: 22 Oct 2018, 13:46
 |  
| Author: | 
Adriano Violante  | 
 | Location: | 
General  | 
 | Severity: | 
Information  | 
 | Subject: | 
Status report after dispatch meeting 09.30  | 
 
  |  
 
 | 
SPB 
FXE 
	- Pulse on demand pump and probe experiment successfully performed.
 
	- The support groups got  the Jungfrau detector working for the users run. 
 
	- VAC sample environment  working.
 
 
SQS 
	- MPS conditions update: Loops will be restarted tomorrow. Tests to then be done in the afternoon and on Wednesday.
 
	- Loop 4 and 5 motors need to be updated . To be discussed with AE when is the best time.
 
	- Pulse on demand capability needed. SQS will contact the CAS contact person.
 
	- Ph. Science camera can be connected again.
 
 
SCS 
	- First shifts succesfull. Support from the support groups is appreciated.
 
	- DAQ groups that have been overlooked are now needed. ITDM asks SCS to send a ticked with the request and the required excel sheet.
 
	-  ITDM also points out, that only DAQ groups that have been already properly commissioned should be used during night shifts.
 
	- Minor problems with gauges. Will be discussed with AE.
 
	- Problems with motors (sometimes they work, sometimes not) has to be discussed in details with AE.
 
 
MID 
  
HED 
XPD 
VAC 
AE 
ITDM 
AoB  |